OpenStudy (anonymous):

Combine like terms. You may use the properties to help: 12v2 − v2 Answer (A) 11v2 (B) 12 (C)12v (D)11v

Parth (parthkohli):

\( \color{Black}{\Rightarrow 12v^2 - v^2 = 12v^2 - 1v^2}\) Agree?

OpenStudy (anonymous):

UM IDK what that means lol

Parth (parthkohli):

Uh-oh. \( \color{Black}{\Rightarrow 1v^2 = v^2}\) right?

OpenStudy (anonymous):

Hmm so whats the answer cause im lost?

OpenStudy (anonymous):

oh i see

Parth (parthkohli):

Well... we can't tell you the answer, as we are NOT allowed to.

OpenStudy (anonymous):

12v2 − v2 is just 12v2 − 1v2

Parth (parthkohli):

http://openstudy.com/code-of-conduct This must be referred.

Parth (parthkohli):

Yep, and then \(12 - 1v^2\) is just \((12 - 1)v^2\).

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so its 11v^2

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

Thanks :) @laxplayer ... and Ok @ParthKohli

OpenStudy (anonymous):

since 12v^2 and 1v^2 is the same u subtract 12- 1 and keep v^2 the same

OpenStudy (anonymous):

Oh i see

Parth (parthkohli):

NO!

Parth (parthkohli):

You are not allowed to give answers, @laxplayer

OpenStudy (anonymous):

oh sorry.......

OpenStudy (anonymous):

@LAGIRL69 ur welcome